OpenStudy (anonymous):

Consider the surface given by 2x^2+5y^2+z^2=4. The surface is sliced by planes of the form x=a, y=b or z=c for values of a, b, c given below. How do I match the resulting slices (in other words, the corresponding traces) with their correct plots??

OpenStudy (anonymous):

-- Also the slices are: 1. z= 1.5 2. x= 1 3. z=0 4. y=0

OpenStudy (anonymous):

Let the value take that on. If you let z be zero we get: 2x^2+5y^2=4 (1/2)x^2+(5/4)y^2=1 which is an ellipse in the x-y plane. They all do something similar. Gotta remember conics.